Please help me with this!!!! Thank you!!!
~~~The length of a rectangle is 5 centimeters less than twice its width. The perimeter of the rectangle is 26 centimeters. What are the dimensions of the rectangle?

Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

The length of a rectangle is 5 centimeters less than twice its width. The perimeter of the rectangle is 26 centimeters. What are the dimensions of the rectangle?
---------------------
Perimeter = 2(L + W)
---
Let width be "W"
The length is 2W-5
------------------
Equation:
26 = 2[(2W-5)+W]
13 = W-5
W = 18 cm (this is the width)
2W-5 = 31 cm (this is the length)
======================================
Cheers,
Stan H.

Answer by Theo(13342)

L = length of the rectangle
W = width of the rectangle
L = 2*W - 5
P = 26 centimeters
perimeter of a rectangle is 2 times the length plus 2 times the width.
P = 2*L + 2*W
since P = 26 and L = 2*W-5 then substitute to get
26 = 2*(2*W-5) + 2*W
26 = 4*W - 10 + 2*W
26 = 6*W - 10
36 = 6*W
W = 6
since L = 2*W - 5, then L = 12 - 5 = 7
so we have
L = 7
W = 6
since P = 2*W + 2*L we have 26 = 2*7 + 2*6 = 14 + 12 = 36 so this checks out.
since we have L = 2*W - 5 we have 7 = 2*6 - 5 = 12 - 5 = 7 so this checks out.
answer is:
length of the rectangle is 7 centimeters
width of the rectangle is 6 centimeters.